Of the 94 participants in a drug trial who were given a new experimental treatment for arthritis, 57 showed improvement. Of the 89 participants given a placebo, 51 showed improvement. 0.05 Significance

1) Assuming independence between the two variables, find the expected frequency for each cell of the table.

Table of expected frequencies:

Improvement. No improvement
Drug ____________. _______________
Placebo__________. _______________

2) Find the value of the x2 test statistic
X2= _____

3) use the given significance level to find the x2 critical value.
The critical value is ____
Answer #1

Following is 2 by 2 table shows the given data:

Improvement No improvement Total
Drug 57 94-57=37 94
Placebo 51 89-51=38 89
Total 108 75 183

Expected frequencies will be calculated as follows:

Following table shows the expected frequencies:

Improvement No improvement
Drug 55.475 38.525
Placebo 52.525 36.475

Following table shows the calculations for chi square test statistics:

O E (O-E)^2/E
57 55.475 0.04192204
51 52.525 0.04427653
37 38.525 0.06036665
38 36.475 0.06375942
Total 0.21032464

Following is the test statistics:

3)

Degree of freedom: df =( number of rows -1)*(number of columns-1) = (2-1)*(2-1)=1

The critical value for 0.05 significance level is 3.841

Since test statistics is less than 3.841 so we fail to reject the null hypothesis.